# Discovery and Expansion of New Domains within Diffusion Models

### v2

@@ Appendix D Detailed Proofs and Geometric Properties
[…]
@@
According to [6] 2.3.3 Bayes’ theorem for Gaussian variables, we know that $q_{\sigma}(\mathbf{x}_{t-1}|\mathbf{x}_{0})$ is also Gaussian, denoted as $\mathcal{N}(\mu_{t-1},\Sigma_{t-1})$ where:
@@
D.2 Proof of Lemma 4.2
Lemma D.2.
Mixtures of spherical Gaussians in $d$ dimensions can be separated provided their centers are separated by more than $d^{\frac{1}{4}}$ distance (i.e, a separation of $\Omega(d^{\frac{1}{4}})$), and even by $\Omega(1)$ separation with more sophisticated algorithms.
@@
According to existing established understanding (Lemma 2.8 from  [7]), for a $d$-dimensional spherical Gaussian of variance 1, all but $\frac{4}{c^{2}}e^{-\frac{c^{2}}{4}}$ fraction of its mass is within the annulus $\sqrt{d-1}-c\leq r\leq\sqrt{d-1}+c$ for any $c>0$, as illustrated in Fig. 5.
@@
More precisely, we note $\mathbf{x}$ is at the North Pole after the coordination rotation with $\mathbf{x}=(\sqrt{(d)}\pm O(1),0,...)$. At the same time, $\mathbf{y}$ is almost on the equator, we can further rotate the coordinate system so that the component of $\mathbf{y}$ that is perpendicular to the axis of the North Pole is in the second coordinate, with $\mathbf{y}=(O(1),\sqrt{(d)}\pm O(1),...)$. Thus we have:
@@ D.3 Geometric Properties
Recall the three geometric properties as below:
@@
For the first observation, when the sample pairs keep approximately the same distance, the direct implication is that those samples are likely to be drawn from some convex region in the high-dimensional space [53]. One typical example is the spherical structure, where every data points exhibit an equal distance from the center.
The second geometric property suggests that the unknown samples could lie on a regular lattice near a low-dimensional manifold or sub-manifold, where the local geometry of the manifold is approximately Euclidean. However, a less evident implication is that for samples drawn from a high-dimensional Gaussian, this property also holds, as detailed in the next section D.4, and illustrated in Fig. 5(c).
The third geometry property implies that the sample points might be isotropic in nature, who are rotationally symmetric around any point in the space. Therefore, any two points drawn from the distribution are equally likely to lie along any direction in the space. This property is also observed for a high-dimensional Gaussian [7], whose covariance matrix is proportional to the identity matrix.
@@
D.4 High-Dimensional Gaussian
@@
We present below several properties of high-dimensional Gaussian from [7], note those are known and established properties, we therefore omit the detailed proofs in this supplement, and ask readers to refer to the original book if interested.
@@
The above lemma is used as the theoretical justification for the proposed empirical search method in [74]. We also adopt the search method using the Gaussian radius for identifying the operational latent space along the denoising chain to perform the OOD sampling.
@@
Appendix E More Deatils about the Latent Sampling Methods
We present here the detailed algorithms for our proposed latent sampling methods, and discuss many other sampling methods that we have tested during experiments.
E.1 Algorithms
While we have described the procedure of our proposed latent sampling method in Sec. 3, the following Algo. 1 and Algo. 2 include details of the overall pipeline and the geometric optimizations, respectively.
Algorithm 1 Domain Discovery and Expansion within DDPMs
  Input: $N$ raw images $\mathbf{x}_{od}$ from the target unseen domain $\mathcal{D}_{od}$, a pre-trained DDPM $p$ for domain $\mathcal{D}_{id}$, target latent step $t$ for sampling ($t$ is a pre-set hyper-parameter discussed in Sec. 4.2).
  Output: images of the unseen target $\mathbf{x}_{od}^{\prime}\in\mathcal{D}_{od}$ // Step 1: get the inverted OOD encodings $\mathbf{x}_{od,t}$
@@ E.1 Algorithms
     Save the OOD latent $\mathbf{x}^{i}_{od,\tau_{S_{inv}}}$ as $\{\mathbf{x}_{od,t}\}^{i=1,...,N}$
  end for// Step 2: get new OOD encodings $\mathbf{x}^{\prime}_{od,t}$
  $\mu_{est}$leftarrow Mean($\mathbf{x}^{1}_{od,t},\mathbf{x}^{2}_{od,t},...,\mathbf{x}^{N}_{od,t}$})
  $\sigma^{2}_{est}$leftarrow Var($\{\mathbf{x}^{1}_{od,t},\mathbf{x}^{2}_{od,t},...,\mathbf{x}^{N}_{od,t}\}$)
  while true do
     $\bar{\mathbf{x}}_{od,t}\sim\mathcal{N}(\mu_{est},\sigma^{2}_{est})$
     $\mathbf{x}_{od,t}\sim\{\mathbf{x}_{od,t}\}^{i=1,...,N}$
     $\mathbf{x}^{\prime}_{od,t}\leftarrow\small{\textsf{Slerp}}(\mathbf{x}_{od,t},\bar{\mathbf{x}}_{od,t})$ // geometric optimization
     if Pass the geometric optimization in Algo. 2 then
        break
     end if
  end while// Step 3: get denoised $\mathbf{x}^{\prime}_{od}\in\mathcal{D}_{od}$
  $\mathbf{x}^{\prime}_{od}\leftarrow p(\mathbf{x}^{\prime}_{od,t},t)$
Algorithm 2 Latent Sampling with Geometric Optimizations
  Input: A sampled OOD latent encoding $x^{\prime}_{od,t}$, $N$ inverted OOD latent $\{\mathbf{x}^{1}_{od,t},\mathbf{x}^{2}_{od,t},...,\mathbf{x}^{N}_{od,t}\}$, distance tolerance $\omega_{d}$, angle tolerance $\omega_{a}$, $N_{ref}$ OOD reference samples
  Output: True or False // Step 1: get pair-wise distance from the inverted OOD encodings $\mathbf{x}_{od,t}$
  $d_{od}\leftarrow 0$
  for $i=1,2,...,n$ do
     $(p,q)\leftarrow\text{RandomInt}(0,N-1)$
     $d_{od}$ += $Euclidien\_distance(x^{p}_{od,t},x^{q}_{od,t})$
  end for
  $d_{od}\leftarrow d_{od}/n$ // Step 2: Geometric optimization based on pre-defined tolerances
  for $i=1,...,N_{ref}$ do
     $d\leftarrow Euclidien\_distance(x^{i}_{od,t},x^{\prime}_{od,t})$
     if $d<d_{od}-\omega_{d}\text{ or }d>d_{od}+\omega_{d}$ then
        return $\text{return }False$
     end if
  end for
  for $i=1,...,N_{ref}$ do
     $(p,q)\leftarrow\text{RandomInt}(0,N_{ref}-1)$
     $\varphi\leftarrow Angle(\vec{x^{\prime}_{od,t}x^{p}_{od,t}},\vec{x^{\prime}_{od,t}x^{q}_{od,t}})$
     if  $\varphi<60-\omega_{a}\text{ or }\varphi>60+\omega_{a}$ then
        return $\text{return }False$
     end if
  end for
  return $\text{return }True$
E.2 Other Sampling Methods
In addition to the main sampling method introduced in the main paper, we have tested many other sampling methods for mining the qualified OOD latent encodings.
We list those sampling methods below as extra information and provide a brief discussion for each.
Approach 1: Estimated Gaussian Sampling. An intuitive way to achieve the OOD sampling based on our analytical understanding from Sec. 2 is to directly fit the latent encodings with a Gaussian distribution and then sample from the estimated Gaussian. However, we note that the high-dimensional Gaussian estimation itself remains as a challenging and complex research topic, especially in a multi-variant case [71, 4]. In general, a reliable estimation of means and variances requires data samples at least 10 times the dimensionality, known as the “rule of ten” [25], which contradicts our few-shot setup. Empirically, we also observe that the synthesis quality with vanilla Gaussian sampling is not very promising. The key reason for this is the gap between the theoretical foundation and practical model training, as also discussed in our main paper.
Approach 2: MCMC Sampling. As an improved statistical method over vanilla Gaussian sampling, we also tested the MCMC sampling technique, which should provide a better and more precise estimation of the distribution based on the inverted latent samples. However, one practical challenge we encountered during the experiments is that MCMC sampling takes an extremely long time for high-dimensional data in this context (i.e., several days using Metropolis-Hastings and Gibbs, which even exceeds the time required to tune the entire model). Therefore, we do not recommend or include this method in our main paper.
Approach 3: Gaussian Sampling w/ Geometric Optimization. A more practical implementation of the Gaussian sampling is to leverage the geometric properties as the domain-specific criteria to further optimize the quality of latent encodings, just as described in Algo. 2. We note this improves over the vanilla Gaussian sampling, but still qualitatively suffers from the mode interference issue.
Appendix F More Details for Generative Experiments
F.1 Background and Evaluation about the Astrophysical Data
Galaxy Data. The images from the GalaxyZoo dataset [58] are observation data of galaxies that belong to one of six categories - elliptical, clockwise spiral, anticlockwise spiral, edge-on, star/don’t know, or merger. The original data format of those galaxy images are also RGB images, thus “somewhat” similar to natural images, but they contain important morphological information to study the galaxies in astronomy.
@@ F.1 Background and Evaluation about the Astrophysical Data
Radiation Data. For the radiation data from [62], the original format is physical quantity instead of RGB images, which correspond to the dust emission.
Dust is a significant component of the interstellar medium in our galaxy, composed of elements such as oxygen, carbon, iron, silicon, and magnesium. Most interstellar dust particles range in size from a few molecules to 0.1 mm (100 $\mu$m), similar to micrometeoroids. The interaction of dust particles with electromagnetic radiation depends on factors like their cross-section, the wavelength of the radiation, and the nature of the grain, including its refractive index and size. The radiation process for an individual grain is defined by its emissivity, which is influenced by the grain’s efficiency factor and includes processes such as extinction, scattering, absorption, and polarization.
In RGB images of dust emission, different colors represent emissions at three wavelengths: blue for 4.5 $\mu$m, green for 24 $\mu$m, and red for 250 $\mu$m. The blue color typically indicates short-wavelength dust emission from point sources, such as young stars or young stellar objects. The green color represents mid-wavelength dust emission from warm and hot dust. The red color signifies long-wavelength dust emission from cold dust.
@@
F.2 More Experimental Results
@@
Figure 6: Illustration of unseen trajectory bandwidth at different diffusion steps. We show qualitative examples using the iDDPM [36] trained on AFHQ-Dog-256 as the base model, the examples of church and cat are both unseen domain images. The image in green boxes indicates the bandwidth we have empirically selected to preserve the reconstruction quality. Compared to the trained image domain (i.e., dogs), cats have a smaller domain gap than churches. Different from the conventional understanding that a smaller domain gap is beneficial for better and easier generalization from a trained model, we observe a larger domain gap signifies a larger bandwidth, making it easier to perform the OOD sampling and synthesis.
F.2.1 Discussion on the Latent Step $t$, Stochasticity and Mode Interference
@@
In our main paper, we briefly discuss the impact of the latent diffusion step $t$ where we perform the inversion and OOD latent sampling. While we empirically find that $t\approx 800$ is a reasonable range for the choice of $t$, we note there exists an entangled mechanism for the trade-off between the sampling difficulty and the mode interference issue.
@@
For the diffusion step $t$, recent studies [74, 64] suggest that $t$ characterizes the formation of image information at different stages of the denoising process. Intuitively, the early stage of the denoising process (e.g., $t>800$) represents a rather chaotic process, the mixing step $t_{m}$ [74] signifies a critical stage where the image semantic information starts to form, and the later stage where $t$ is close to 0 demonstrates a stage during which more fine-grained pixel-level information are introduced to the final generated data. From the distribution point of view, the influence of $t$ can be interpreted as the convergence of distributions, where $t=T$ is a standard Gaussian by definition, thus the ID and OOD modes are more difficult to separate. However, as the denoising process gets closer to the real image space at $t=0$, the sampling difficulty increases as the implicit distribution moves away from the standard Gaussian.
@@
Meanwhile, the diffusion step $t$ is not the only factor that impacts the trade-off between sampling difficulty and mode interference. While scarcely discussed in the main paper, we note the stochasticity of the denoising trajectory also plays a similar role as the diffusion step in this work. The stochasticity of the denoising trajectory in DMs has been proven to be generally beneficial in improving the synthesis quality [27, 29, 31, 74]. In this work, while we choose the $\eta=0$ for the main paper, a tolerance for a certain range of stochasticity allows us to follow a “relatively deterministic” denoising process $p_{\eta=k}$, with $k\neq 0$, instead of the completely deterministic $p_{i}$. We hereby refer to it as “bandwidth of the unseen trajectories,” denoted as $\mathcal{B}_{\eta,t}$, which can be used to quantify the “mode interference”. Another interpretation is to analog the trajectory bandwidth $\mathcal{B}_{\eta,t}$ to the actual subspace volume occupied by the OOD latent samples. Fig. 6 shows more qualitative results for the bandwidth search in the reconstruction task and reveals its connection to the diffusion step $t$. Overall, the bandwidth is a hyper-parameter that relates to the base model and the unseen domains, and the diffusion step $t$, while the bandwidth gets larger at the latent spaces closer to the raw image domains, sampling from OOD unseen distributions also gets more difficult.
F.2.2 Discussion on Model Designs

### v3

@@ C.3 Parameter-Independent Properties: Gaussian Priors
As also mentioned in [song2020ddim], one can derive Lemma 4 by assuming for any $t\leq T$, $q_{\sigma}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I})$ holds, if:
and then prove the statement with an induction argument for $t$ from $T$ to 1, since the base case $(t=T)$ already holds by definition. We provide the detailed proof below.
@@
$q_{\sigma}(\mathbf{x}_{t-1}|\mathbf{x}_{0}):=\int_{\mathbf{x}_{t}}q_{\sigma}(\mathbf{x}_{t}|\mathbf{x}_{0})q_{\sigma}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})d\mathbf{x}_{t},$ (7)
@@
According to [bishop2006pattern] 2.3.3 Bayes’ theorem for Gaussian variables, we know that $q_{\sigma}(\mathbf{x}_{t-1}|\mathbf{x}_{0})$ is also Gaussian, denoted as $\mathcal{N}(\mu_{t-1},\Sigma_{t-1})$ where:
@@
C.4 Data-Dependent Properties: Mode Interference and Separability
In the literature of GANs-based generative models [gan], “mode collapse” is a common issue that describes the training failure when generated images tend to be very similar given randomly sampled starting encodings from the Gaussian prior. Within the context of diffusion models in our work, we explicitly reveal a phenomenon analog to the “mode collapse” in GANs, which we refer to as “mode interference”, as qualitatively illustrated in Fig. 3 (a).
Notably, a key precondition to achieving the effective OOD latent sampling is that the established OOD prior mode should be separable from the ID Gaussian prior mode (i.e., a standard Gaussian). Otherwise, the denoised image would fall into the training domain as in Fig. 3 (b). The separability is further supported and validated by our empirical verification below in Sec. C.5.
C.5 Analytical Experiments
We show empirical verification from multiple perspectives to support our parameter-independent and data-dependent properties described above.
Geometrical Properties of Gaussians. We leverage the geometrical measurements established of the high-dimensional studies in mathematics [blum2020foundations], as additional empirical support for the Gaussian priors in Sec. C.3. Specifically, we compute several geometric metrics, including the pair-wise angles (angles formed by three arbitrary samples), sample-to-origin angles (angles formed by two arbitrary samples and the origin), pair-wise distance (euclidean distance between two arbitrary samples) and distance between OOD and ID Gaussian centers, and list the results in Tab. 3.
Table 3: Geometric properties of inverted ID and OOD latent encodings. The results are computed based on 1K sample pairs. We report the mean and std for each geometric measurement to ensure the statistical significance. The base model is trained on AFHQ-Dog [choi2020stargan] in 256x256.
[…]
Center-Distance - 33.3 26.0 33.2 54.7 60.8
Clf. Acc. - 0.97 0.99 0.99 1.0 1.0
Characteristics above are typical geometric properties possessed by isotropic high-dimensional Gaussians [blum2020foundations]. Notably, three randomly sampled points from a high-dimensional Gaussian are almost surely form an equilateral triangle and are almost surely nearly orthogonal, which corresponds to the constant 60 pair-wise angle and 90 sample-to-origin angle in the first and second rows of Tab. 3, respectively.
Mode Separability. As revealed by our analysis in Sec. C.4, the separability between ID and OOD Gaussian modes is critical for synthesizing target unseen domain images without modifying the model parameters and for avoiding the “mode interference”. We further provide validation from the statistical and learning-based classifier perspectives to support the separability claim.
Statistical Validation. The separability of high-dimensional Gaussians follows Lemma C.2 [blum2020foundations], which states that spherical Gaussians can be relaxingly separated by $\Omega(d^{\frac{1}{4}})$, or even $\Omega(1)$ with more sophisticated algorithms. In other words, for a DDPM trained on $\times 256$ images with dimensionality $d=3\times 256\times 256$, ID and OOD modes can be well separated and avoid interference given a distance larger than $d^{\frac{1}{4}}\approx 21$, which is further validated by the empirical distance between centers, listed in the forth row of Tab. 3.
Lemma C.2
Mixtures of spherical Gaussians in $d$ dimensions can be separated provided their centers are separated by more than $d^{\frac{1}{4}}$ distance (i.e., a separation of $\Omega(d^{\frac{1}{4}})$). and even by $\Omega(1)$ separation with more sophisticated algorithms.
@@ C.5 Analytical Experiments
According to existing established understanding (Lemma 2.8 from  [blum2020foundations]), for a $d$-dimensional spherical Gaussian of variance 1, all but $\frac{4}{c^{2}}e^{-\frac{c^{2}}{4}}$ fraction of its mass is within the annulus $\sqrt{d-1}-c\leq r\leq\sqrt{d-1}+c$ for any $c>0$, as illustrated in Fig. 4.
@@
[…]
@@
Classifier Validation. Another empirical perspective to validate the separability between modes in the latent spaces is using the classifiers as in existing literature [shen2020interpreting, zhu2023boundary]. Specifically, a linear classifier such as SVMs [hearst1998support] can be fitted to test the separability between ID and OOD encodings in the latent spaces. In our analytical experiments, we fit SVMs on 1K inverted ID and OOD samples following the 7:3 training-testing ratio, and report the test accuracy in Tab. 3. As additional clarification, the classification results are obtained with the test on the latent space $\mathcal{X}_{T}$. Our rationale behind the choice of $T$ corresponds to the recent findings of DMs [zhu2023boundary, yang2023cow], which indicates that $\mathcal{X}_{T}$, as the departure latent space, has the largest probabilistic support for the trained domain. In other words, if the latent ID and OOD modes can be separated in $\mathcal{X}_{T}$, they can be separated more easily in other $\mathcal{X}_{t}$, for $t=\{T-1,...,t,...1\}$.
@@ C.6 Geometric Properties
The three geometric properties as below:
@@
For the first observation, when the sample pairs keep approximately the same distance, the direct implication is that those samples are likely to be drawn from some convex region in the high-dimensional space [wang2012geometric]. One typical example is the spherical structure, where every data points exhibit an equal distance from the center.
The second geometric property suggests that the unknown samples could lie on a regular lattice near a low-dimensional manifold or sub-manifold, where the local geometry of the manifold is approximately Euclidean. However, a less evident implication is that for samples drawn from a high-dimensional Gaussian, this property also holds, as detailed in the next section C.7, and illustrated in Fig. 4(c).
The third geometry property implies that the sample points might be isotropic in nature, who are rotationally symmetric around any point in the space. Therefore, any two points drawn from the distribution are equally likely to lie along any direction in the space. This property is also observed for a high-dimensional Gaussian [blum2020foundations], whose covariance matrix is proportional to the identity matrix.
@@
C.7 High-Dimensional Gaussian
@@
We present below several properties of high-dimensional Gaussian from [blum2020foundations], note those are known and established properties, we therefore omit the detailed proofs in this supplement, and ask readers to refer to the original book if interested.
@@
The above lemma is used as the theoretical justification for the proposed empirical search method in [zhu2023boundary]. We also adopt the search method using the Gaussian radius for identifying the operational latent space along the denoising chain to perform the OOD sampling.
@@
Appendix D More Details about the Latent Sampling Methods
Intuitively, unlike sampling from a pure Gaussian, sampling from an approximate-Gaussian in non-trivial. Alternatively, the general property of Gaussian distributions ensures that any convex combination of two samples lies within the same distribution. This motivates our choice of a simple yet effective approach to identify additional qualified $\mathbf{x}^{\prime}_{ood,T}\in\mathcal{X}_{ood,T}$ using latent interpolation, a technique widely adopted in generative modeling for geometric and spatial analysis [song2020ddim, shen2020interpreting, zhu2023boundary, baumann2024continuous, preechakul2022diffusion]. Specifically, as shown in Algo. 1, new latent samples $\mathbf{x}^{\prime}_{od,T}$ are obtained by interpolating between two known latent encodings $\mathbf{x}^{i}_{od,T}$ and $\mathbf{x}^{j}_{od,T}$, both inverted from raw data.
Algorithm 1 Latent sampling-centric approach for domain generalization
Input: Arbitrary pre-trained DDPM $p_{\theta}$ in $\mathcal{D}_{id}$, NN images $\mathbf{x}_{od}\in\mathcal{D}_{od}$.
Output: images of the unseen target domain $\mathbf{x}_{od}^{\prime}\in\mathcal{D}_{od}$ // Get the inverted latent encodings $\mathbf{x}_{od,T}$
@@ Appendix D More Details about the Latent Sampling Methods
end for// Get new ood encodings $\mathbf{x}^{\prime}_{od,T}$ and denoise via DDIMs
$\mathbf{x}^{\prime}_{od,T}\leftarrow\text{Interp}(\mathbf{x}^{i}_{od,T},\mathbf{x}^{j}_{od,T}),\text{for}\>i\neq j,\>(i,j)\in\{1,2,...,N\}$
$\mathbf{x}^{\prime}_{od}\leftarrow p_{\theta}(\mathbf{x}^{\prime}_{od,T},T)$
Appendix E More Details for Generative Experiments
E.1 Background and Evaluation about the Astrophysical Data
Galaxy Data. The images from the GalaxyZoo dataset [willett2013galaxy] are observation data of galaxies that belong to one of six categories - elliptical, clockwise spiral, anticlockwise spiral, edge-on, star/don’t know, or merger. The original data format of those galaxy images are also RGB images, thus “somewhat” similar to natural images, but they contain important morphological information to study the galaxies in astronomy.
@@ E.1 Background and Evaluation about the Astrophysical Data
Radiation Data. For the radiation data from [xu2023predicting], the original format is physical quantity instead of RGB images, which correspond to the dust emission.
[…]
@@
Qualitative evaluation from astrophysicists. As the evaluation of astrophysical data requires deep domain expertise, we collaborated with astrophysicists to subjectively assess the quality of the generated data using Mean Opinion Scores (MOS). Specifically, we provided 50 non-cherry-picked generated samples from our unseen domain generalization experiments alongside 50 raw data samples from true physical simulations. For each generated sample across the two astrophysical datasets, a score ranging from 1 to 5 was assigned relative to the raw data, where 5 indicates the highest quality and 1 is the lowest.
[…]
E.2 More Experimental Results
@@
Figure 5: Illustration of unseen trajectory bandwidth at different diffusion steps. We show qualitative examples using the iDDPM [nichol2021improveddmp] trained on AFHQ-Dog-256 as the base model, the examples of church and cat are both unseen domain images. The image in green boxes indicates the bandwidth we have empirically selected to preserve the reconstruction quality. Compared to the trained image domain (i.e., dogs), cats have a smaller domain gap than churches. Different from the conventional understanding that a smaller domain gap is beneficial for better and easier generalization from a trained model, we observe a larger domain gap signifies a larger bandwidth, making it easier to perform the OOD sampling and synthesis.
E.2.1 Discussion on the Latent Step $t$, Stochasticity and Mode Interference
@@
While we empirically find that $t\approx 800$ is a reasonable range for the choice of $t$, we note there exists an entangled mechanism for the trade-off between the sampling difficulty and the mode interference issue.
@@
For the diffusion step $t$, recent studies [zhu2023boundary, yang2023cow] suggest that $t$ characterizes the formation of image information at different stages of the denoising process. Intuitively, the early stage of the denoising process (e.g., $t>800$) represents a rather chaotic process, the mixing step $t_{m}$ [zhu2023boundary] signifies a critical stage where the image semantic information starts to form, and the later stage where $t$ is close to 0 demonstrates a stage during which more fine-grained pixel-level information are introduced to the final generated data. From the distribution point of view, the influence of $t$ can be interpreted as the convergence of distributions, where $t=T$ is a standard Gaussian by definition, thus the ID and OOD modes are more difficult to separate. However, as the denoising process gets closer to the real image space at $t=0$, the sampling difficulty increases as the implicit distribution moves away from the standard Gaussian.
@@
Meanwhile, the diffusion step $t$ is not the only factor that impacts the trade-off between sampling difficulty and mode interference. While scarcely discussed in the main paper, we note the stochasticity of the denoising trajectory also plays a similar role as the diffusion step in this work. The stochasticity of the denoising trajectory in DMs has been proven to be generally beneficial in improving the synthesis quality [karras2022elucidating, kim2022diffusionclip, kwon2022diffusion, zhu2023boundary]. In this work, while we choose the $\eta=0$ for the main paper, a tolerance for a certain range of stochasticity allows us to follow a “relatively deterministic” denoising process $p_{\eta=k}$, with $k\neq 0$, instead of the completely deterministic $p_{i}$. We hereby refer to it as “bandwidth of the unseen trajectories,” denoted as $\mathcal{B}_{\eta,t}$, which can be used to quantify the “mode interference”. Another interpretation is to analog the trajectory bandwidth $\mathcal{B}_{\eta,t}$ to the actual subspace volume occupied by the OOD latent samples. Fig. 5 shows more qualitative results for the bandwidth search in the reconstruction task and reveals its connection to the diffusion step $t$. Overall, the bandwidth is a hyper-parameter that relates to the base model and the unseen domains, and the diffusion step $t$, while the bandwidth gets larger at the latent spaces closer to the raw image domains, sampling from OOD unseen distributions also gets more difficult.
E.2.2 Discussion on Model Designs